OpenStudy (anonymous):

yes you may I need help with write an equation in slope intercept form for the line that passes through the given points (2,3), (1,5)

OpenStudy (anonymous):

First you must calculate the slope between these two points....that works out to be -2 in this case. Now use the slope and one of the points in the equation y-y=m(x-x) so using the point (1,5) you have y-5=-2(x-1) and solve this for y and you will have the slope intercept form.....